OpenStudy (anonymous):

Molly is having her birthday party at Pizza Party Palace. The costs involved are listed below. Pizza Party Palace requires a 15% deposit prior to the date of the event. If the deposit has been paid, how much is still owed for Molly's party? Party Prices • Food & Drink $200 • Game Tokens $150 • Other $70

OpenStudy (jdoe0001):

\(\large { \begin{array}{lrlll} prices \\\hline\\ food\ \& \ Drink&200\\ Game \ Tokens& 150\\ other& +\ 70\\ \hline\\ &\square \end{array}\quad 15\%\ of \ \square \to \frac{15}{100}\cdot \square }\) how much money is owed still? well... find the 15% of the total, and subtract it from the total 15% has been paid the rest is what is owed